DIFFERENTIAL GEOMETRY/PDE SEMINAR

Friday, Sept. 14, 2001 Padelford C-401 2:00 pm

Constructing invariants of isolated hypersurface singularities

Michael G. Eastwood (University of Adelaide and MSRI)

John Mather and Stephen Yau showed that isolated singularities of complex hypersurfaces are classified by their associated moduli algebras. This suggests that one should be able to associate invariants to a hypersurface singularity via its moduli algebra. I shall explain how this can be done.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, https://www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.